A particle oscillates along a X-axis according to equation x =0.05 sin (5t – pi/6). Find velocity at t=0 second.
 

Grade:12th pass

2 Answers

Arun
6 years ago
Dear student
 
For velocity
V = d x/dt = 0.05 cos(5t- pi/6) * 5
Hence at t = 0
V = 0.25 cos (-pi/6)
V = 0.25 *(1/2) *\sqrt3
 
Regards
Arun (askIITians forum expert)
Ayushi Singh
one year ago
As we know that
V= dx /dt = d 0.05 Sin (5t -π/6 )/dt
V = 0.05 ×5 Cos (5t -π/6)
V = 0.05 ×5 Sin (5t -π/6 +π/2 )
V = 0.25 Sin (5×0 + π/3 )
V = 0.25 Sin π/3 
V = 0.25 √3/2
V=  √3/8 m/ sec